# Hovering Flight in Flapping Insects and Hummingbirds: A Natural Real-Time and Stable Extremum Seeking Feedback System

### v4

@@ Abstract
In this paper, we take an initial and novel step toward characterizing the hovering phenomenon in flapping insects and hummingbirds as a new class of extremum seeking (ES) feedback systems. By characterizing hovering flight in insects and hummingbirds as a natural hovering ES system, we achieve: (1) very simple, (2) stable, (3) model-free, and (4) real-time hovering. More importantly, our hovering ES characterization only needs the natural oscillations of the wing as the ES input. That is, unlike other control techniques in the literature, the natural hovering ES system only needs the natural flapping action built in the system, and feedback of local sensations (measurements) related to the altitude where the insect seeks to stabilize itself. Said ES characterization, can become an important initial step in starting a new line of research that may succeed in resolving the long-standing gap between model-based control theory and the biologically observed mechanisms that stabilize hovering flight. We provide simulation trials, including comparisons with some approaches from literature, to demonstrate the effectiveness and robustness of our results. We used literature data for hawkmoth, cranefly, bumblebee, dragonfly, hoverfly, and a hummingbird.
@@ I Introduction
The long-held view and consensus in the literature is that hovering flapping flight is inherently unstable in the open-loop sense and active control may be required for stabilization [1, 9, 12, 13, 14, 15, 16, 17]. A recent study [18] extended stability analysis to a broader range of hovering insects has reached the same conclusion as well. Recently, the authors in [19, 20] argued that concluding instability is due to a deficiency in first-order averaging and that hovering flight may be shown stable—including the pitching motion in an open-loop sense if higher-order averaging is used; however, they acknowledged that feedback control may be required for other insects to stabilize their hovering flight. Natural and biological observations/experiments are in consensus that flapping insects rely heavily on their sensation, which means it is highly plausible that they use some feedback control mechanism to stabilize their hovering in real time [21, 22, 12, 17, 23, 24]. The authors in [17, 22] examined how freely flying fruit flies react in real time to external disturbances affecting their yaw and pitch: small ferromagnetic materials were attached to the flies, allowing for reorientation via a magnetic field, and observations indicated that the flies adapted their wing kinematics in real-time. Thus, this can imply that, in the absence of passive damping, active sensorimotor feedback is important for stabilization.
@@
From a control-theoretic perspective, stabilized hovering is achieved through sensory feedback [20, 25, 12, 17, 26, 21]. This understanding has motivated many researchers to develop feedback-based control techniques aimed at enabling stable hovering. Consequently, a variety of model-dependent control strategies have been proposed in the literature [1, 9]. Standard approaches such as Linear Quadratic Regulator (LQR), Proportional-Integral-Derivative (PID) control, and other nonlinear designs rely on dynamic and aerodynamic models that require knowledge of the equilibrium hovering point a priori [27, 28, 29, 30]. In several studies, aerodynamic models necessary for control were based on computational fluid dynamics simulations, which are inherently non-real-time [31, 28, 32, 33]. It is worth mentioning that, although PID and its variants (PI and PD) are real-time controllers, even the PI scheme that reproduces yaw recovery in fruit flies [17] — as the authors themselves state — requires prior estimation of morphological and aerodynamic parameters and subsequent hand-tuning of gain parameters, with all parameters varying by ±$\pm 15\%$. Moreover, none of the control methods, techniques, or mechanisms discussed above leverage the inherent high-frequency wing-flapping perturbation action, which can serve as a built-in probing signal.
@@
Motivation. The consensus that hovering flight is open-loop unstable and needs feedback control for stabilization has motivated researchers to introduce a variety of feedback control techniques to achieve stable hovering; however, available control methods can be model-dependent, require knowledge of the equilibrium hovering point a priori, and can potentially be computationally expensive to implement in real-time due to tuning challenges and/or high need of computational power. Moreover, none of these controllers leverage the inherent high-frequency wing-flapping perturbation action built into the system. Hence, we are motivated to propose a control method that can frame the hovering problem while using the built-in high-frequency wing perturbation action in a natural way as the probing signal for the controller itself. For that, we turn our attention to Extremum Seeking (ES) control systems [34, 35, 36], which utilize perturbation actions as the probing signal for the controller to derive a given dynamical system towards the extremum of unknown objective functions in a model-free, real-time, sensation-based fashion. Moreover, ES feedback control systems have achieved breakthroughs in characterizing and mimicking challenging behaviors in biological organisms immersed in fluid structures — see, for example, fish [37] and microswimmers [38, 39]. Recently, our group introduced ES characterization in flying organisms and succeeded in decoding the energy-efficient soaring strategy of birds [40, 41].
@@
Contribution. In this paper, we take an initial and novel step toward characterizing the hovering phenomenon in flapping insects and hummingbirds as a new class of ES feedback systems. By characterizing hovering flight in insects and hummingbirds as a natural hovering ES system, we achieve: (1) very simple, (2) stable, (3) model-free, and (4) real-time hovering. More importantly, our hovering ES characterization only needs the natural oscillations of the wing (i.e., the natural flapping action, built in the system, unlike other control techniques in the literature) as the ES input, and feedback of local sensations (measurements) related to the altitude where the insect seeks to stabilize itself. Said ES characterization, can become an important initial step in starting a new line of research that may succeed in resolving the long-standing gap between model-based control theory and the biologically observed mechanisms that stabilize hovering flight. It also may provide model-free, real-time, computationally cheap, bio-mimicry possibilities for flapping robots.
@@ II Modelling of Flapping Insects and Hummingbirds: Body Dynamics and Aerodynamics
[…]
@@ III.2 Re-examining Some Literature Results from the Eyes of Natural Hovering Extremum Seeking
Figure 6: Open-loop simulation (red) vs. closed-loop simulation (black). Open-loop velocity trajectory is stabilizes about very small non-zero velocity equilibrium due to numerical errors, leading to slow linear descending in altitude. This is not the case with ES closed-loop feedback, which achieves successfully, and able to maintain, stable hovering.
@@ III.3 Comparison Between PID and Natural Hovering Extremum Seeking Approach
Figure 7: Comparison of altitude stabilization using PID (red) vs. natural hovering ES system (black). Natural hovering ES system is performing better in terms of settling time and its lesser overshoot.
@@
$-0.10$ $2.411\mathrm{e}{-02}$ $1.027\mathrm{e}{-01}$ $2.296\mathrm{e}{-02}$
$-0.20$ $2.386\mathrm{e}{-02}$ $1.152\mathrm{e}{-01}$ $6.272\mathrm{e}{-03}$
$-0.30$ $2.188\mathrm{e}{-02}$ $8.514\mathrm{e}{-02}$ $4.248\mathrm{e}{-03}$
@@
Figure 8: Comparison of altitude stabilization using the optimal PID controller (red) versus the natural hovering ES system (black). The ES loop settles faster and with smaller overshoot for both initial conditions ($w_{0}=0.2$ and $w_{0}=-0.2$), while using the same ESC feedback gain in every case and different optimal PID gains as shown in table 7 .
@@ IV Stability analysis of our natural hovering extremum seeking system and comparison with open-loop
where $[\bm{Y},\bm{Z}]$ is the Lie bracket between the vector fields $\bm{Y}$ and $\bm{Z}$, defined as
$[\bm{Y},\bm{Z}]=\frac{\partial\bm{Z}}{\partial\bm{x}}\bm{Y}-\frac{\partial\bm{Y}}{\partial\bm{x}}\bm{Z}.$ (15)
@@
Hawkmoth $-5.11e{+05}$ $-4.81e{+05}$ $-5.25e{+05}$ $-5.48e{+05}$
$-1.78e{+04}$ $-1.66e{+04}$ $-1.82e{+04}$ $-1.90e{+04}$
Cranefly $-1.50e{+06}$ $-1.60e{+06}$ $-1.60e{+06}$ $-1.68e{+06}$
$-9.28e{+04}$ $-1.01e{+05}$ $-9.60e{+04}$ $-1.04e{+05}$
Bumblebee $-1.20e{+07}$ $-1.31e{+07}$ $-1.26e{+07}$ $-1.27e{+07}$
$-1.02e{+05}$ $-1.12e{+05}$ $-1.07e{+05}$ $-1.08e{+05}$
Dragonfly $-4.14e{+06}$ $-4.50e{+06}$ $-4.39e{+06}$ $-4.52e{+06}$
$-1.41e{+05}$ $-1.54e{+05}$ $-1.60e{+05}$ $-1.82e{+05}$
Hoverfly $-7.81e{+06}$ $-8.55e{+06}$ $-8.04e{+06}$ $-8.11e{+06}$
$-1.20e{+05}$ $-1.31e{+05}$ $-1.24e{+05}$ $-1.25e{+05}$
Hummingbird $-3.33e{+06}$ $-3.65e{+06}$ $-4.01e{+06}$ $-4.38e{+06}$
$-2.00e{+05}$ $-2.14e{+05}$ $-2.40e{+05}$ $-2.62e{+05}$
@@
Additionally, we made computations and included also in Table 10 the eigenvalues of stable open-loop hovering [46]; the reader is encouraged to recall relevant comments/discussions about open-loop stability provided in Subsection III.A. To obtain the optimal–PID eigenvalues, the averaged model was first derived exactly as in the open-loop case. Classical linear-control theory was then applied to incorporate the PID loop and evaluate the closed-loop eigenvalues, following the procedure in [59]. All computational steps are available in the accompanying code repository in [60]. It is important to mention that there is a difference in the nature of the stability between PID/ES on one hand and the open-loop case on the other. In the open-loop configuration, the equilibrium need not correspond to hovering ($w=0$). Stability can occur at a non-zero $w$, leading to steady ascent or descent, as discussed in Section III.B; only by selecting the exact trim (numerically demanding) does one obtain open-loop hovering. By contrast, both the PID loop and the ES controller always stabilize to $w\!=\!0$, so any disturbance is rejected and altitude is recovered. It can also be noted that in some instances the optimal PID closed-loop is less effective than the open-loop system—in the hawkmoth case, for example, as shown in Table 10 (but this should be put in the context that PID stabilization is hovering stabilization, which is not the case in the open-loop case. Overall, the eigenvalue results confirm the improvement in hovering stability by the natural hovering ES system using different objective functions compared to the open-loop case and optimal PID case. The results confirm that via extremum seeking control (ESC), hovering exhibits eigenvalues with more negative real components (i.e., shifts the eigenvalues further into the left-half complex plane), indicating stronger stability properties and more robustness.

### v6

@@ Abstract
In this paper, we take an initial and novel step toward characterizing the physics of the hovering phenomenon in flapping insects and hummingbirds as a new class of extremum seeking (ES) feedback systems. By characterizing hovering flight in insects and hummingbirds as a natural hovering ES system, we achieve: (1) very simple, (2) stable, (3) model-free, and (4) real-time hovering. More importantly, our hovering ES characterization only needs the natural oscillations of the wing as the ES input. That is, unlike other control techniques in the literature, the natural hovering ES system only needs the natural flapping action built in the system, and feedback of local sensations (measurements) related to the altitude where the insect seeks to stabilize itself. Said ES characterization, can become an important initial step in starting a new line of research that may succeed in resolving the long-standing gap between model-based control theory and the biologically observed mechanisms that stabilize hovering flight. We provide simulation trials, including comparisons with some approaches from literature, to demonstrate the effectiveness and robustness of our results. We used literature data for hawkmoth, cranefly, bumblebee, dragonfly, hoverfly, and a hummingbird.
@@ I Introduction
The long-held view and consensus in the literature is that hovering flapping flight is inherently unstable in the open-loop sense and active control may be required for stabilization [1, 9, 12, 13, 14, 15, 16, 17]. A recent study [18] extended stability analysis to a broader range of hovering insects has reached the same conclusion as well. Recently, the authors in [19, 20] argued that concluding instability is due to a deficiency in first-order averaging and that hovering flight may be shown stable—including the pitching motion in an open-loop sense if higher-order averaging is used; however, they acknowledged that feedback control may be required for other insects to stabilize their hovering flight. Natural and biological observations/experiments are in consensus that flapping insects rely heavily on their sensation, which means it is highly plausible that they use some feedback control mechanism to stabilize their hovering in real time [21, 22, 12, 17, 23, 24].
@@
From a control-theoretic perspective, stabilized hovering is achieved through sensory feedback [20, 25, 12, 17, 26, 21]. This understanding has motivated many researchers to develop feedback-based control techniques aimed at enabling stable hovering. Consequently, a variety of model-dependent control strategies have been proposed in the literature [1, 9]. Standard approaches such as Linear Quadratic Regulator (LQR), Proportional-Integral-Derivative (PID) control, and other nonlinear designs rely on dynamic and aerodynamic models that require knowledge of the equilibrium hovering point a priori [27, 28, 29, 30]. In several studies, aerodynamic models necessary for control were based on computational fluid dynamics simulations, which are inherently non-real-time [31, 28, 32, 33]. It is worth mentioning that, although PID and its variants (PI and PD) are real-time controllers, even the PI scheme that reproduces yaw recovery in fruit flies [17] — as the authors themselves state — requires prior estimation of morphological and aerodynamic parameters and subsequent hand-tuning of gain parameters, with all parameters varying by $\pm 15\%$. Moreover, none of the control methods, techniques, or mechanisms discussed above leverage the inherent high-frequency wing-flapping perturbation action, which can serve as a built-in probing signal.
@@
Motivation. The consensus that hovering flight is open-loop unstable and needs feedback control for stabilization has motivated researchers to introduce a variety of feedback control techniques to achieve stable hovering; however, available control methods can be model-dependent, require knowledge of the equilibrium hovering point a priori, and can potentially be computationally expensive to implement in real-time due to tuning challenges and/or high need of computational power. Moreover, none of these controllers leverage the inherent high-frequency wing-flapping perturbation action built into the system. Hence, we are motivated to propose a control method that can frame the hovering problem while using the built-in high-frequency wing perturbation action in a natural way as the probing signal for the controller itself. For that, we turn our attention to Extremum Seeking (ES) control systems [34, 35, 36], which utilize perturbation actions as the probing signal for the controller to derive a given dynamical system towards the extremum of unknown objective functions in a model-free, real-time, sensation-based fashion. Moreover, ES feedback control systems have achieved significant successes in characterizing and mimicking challenging behaviors in biological organisms immersed in fluid structures — see, for example, fish [37] and microswimmers [38, 39]. Recently, our group introduced ES characterization in flying organisms and succeeded in decoding the energy-efficient soaring strategy of birds [40, 41].
@@
Contribution. In this paper, we take an initial and novel step toward characterizing the physics of the hovering phenomenon in flapping insects and hummingbirds as a new class of ES feedback systems. By characterizing hovering flight in insects and hummingbirds as a natural hovering ES system, we achieve: (1) very simple, (2) stable, (3) model-free, and (4) real-time hovering. More importantly, our hovering ES characterization only needs the natural oscillations of the wing (i.e., the natural flapping action, built in the system, unlike other control techniques in the literature) as the ES input, and feedback of local sensations (measurements) related to the altitude where the insect seeks to stabilize itself. Said ES characterization, can become an important initial step in starting a new line of research that may succeed in resolving the long-standing gap between model-based control theory and the biologically observed mechanisms that stabilize hovering flight. It also may provide model-free, real-time, computationally cheap, bio-mimicry possibilities for flapping robots.
@@ II Modelling of Flapping Insects and Hummingbirds: Body Dynamics and Aerodynamics
where $\rho$ represents the air density, $\alpha_{m}$ is the mean angle of attack sustained throughout the flapping hovering. The total mass of the insect, including both wing and body mass, is represented by $m$, and the constants $I_{mn}$ are calculated based on the chord distribution $c(r)$ along the wing, given in [50] by the integral $I_{mn}=2\int_{0}^{R}r^{m}c^{n}(r)dr$. Note that $C_{L_{\alpha}}$ is the lift curve slope of the wing as defined in [45]:
@@ III.2 Re-examining Some Literature Results from the Eyes of Natural Hovering Extremum Seeking
Figure 6: Open-loop simulation (red) vs. closed-loop simulation (blue). Open-loop velocity trajectory is stabilizes about very small non-zero velocity equilibrium due to numerical errors, leading to slow linear descending in altitude. This is not the case with ES closed-loop feedback, which achieves successfully, and able to maintain, stable hovering.
@@ III.3 Comparison Between PID and Natural Hovering Extremum Seeking Approach
Figure 7: Comparison of altitude stabilization using PID (red) vs. natural hovering ES system (blue). Natural hovering ES system is performing better in terms of settling time and its lesser overshoot.
@@
0.10-0.10 $2.411\mathrm{e}{-02}$ $1.027\mathrm{e}{-01}$ $2.296\mathrm{e}{-02}$
0.20-0.20 $2.386\mathrm{e}{-02}$ $1.152\mathrm{e}{-01}$ $6.272\mathrm{e}{-03}$
0.30-0.30 $2.188\mathrm{e}{-02}$ $8.514\mathrm{e}{-02}$ $4.248\mathrm{e}{-03}$
@@
Figure 8: Comparison of altitude stabilization using the optimal PID controller (red) versus the natural hovering ES system (blue). The ES loop settles faster and with smaller overshoot for both initial conditions ($w_{0}=0.2$ and $w_{0}=-0.2$), while using the same ESC feedback gain in every case and different optimal PID gains as shown in table 7 .
@@ IV Stability analysis of our natural hovering extremum seeking system and comparison with open-loop
where $\bm{Y},\bm{Z}]$ is the Lie bracket between the vector fields $\bm{Y}$ and $\bm{Z}$, defined as
$\bm{Y},\bm{Z}]=\frac{\partial\bm{Z}}{\partial\bm{x}}\bm{Y}-\frac{\partial\bm{Y}}{\partial\bm{x}}\bm{Z}.$ (15)
@@
Hawkmoth 5.11e+05-5.11e{+05} 4.81e+05-4.81e{+05} 5.25e+05-5.25e{+05} 5.48e+05-5.48e{+05}
1.78e+04-1.78e{+04} 1.66e+04-1.66e{+04} 1.82e+04-1.82e{+04} 1.90e+04-1.90e{+04}
Cranefly 1.50e+06-1.50e{+06} 1.60e+06-1.60e{+06} 1.60e+06-1.60e{+06} 1.68e+06-1.68e{+06}
9.28e+04-9.28e{+04} 1.01e+05-1.01e{+05} 9.60e+04-9.60e{+04} 1.04e+05-1.04e{+05}
Bumblebee 1.20e+07-1.20e{+07} 1.31e+07-1.31e{+07} 1.26e+07-1.26e{+07} 1.27e+07-1.27e{+07}
1.02e+05-1.02e{+05} 1.12e+05-1.12e{+05} 1.07e+05-1.07e{+05} 1.08e+05-1.08e{+05}
Dragonfly 4.14e+06-4.14e{+06} 4.50e+06-4.50e{+06} 4.39e+06-4.39e{+06} 4.52e+06-4.52e{+06}
1.41e+05-1.41e{+05} 1.54e+05-1.54e{+05} 1.60e+05-1.60e{+05} 1.82e+05-1.82e{+05}
Hoverfly 7.81e+06-7.81e{+06} 8.55e+06-8.55e{+06} 8.04e+06-8.04e{+06} 8.11e+06-8.11e{+06}
1.20e+05-1.20e{+05} 1.31e+05-1.31e{+05} 1.24e+05-1.24e{+05} 1.25e+05-1.25e{+05}
Hummingbird 3.33e+06-3.33e{+06} 3.65e+06-3.65e{+06} 4.01e+06-4.01e{+06} 4.38e+06-4.38e{+06}
2.00e+05-2.00e{+05} 2.14e+05-2.14e{+05} 2.40e+05-2.40e{+05} 2.62e+05-2.62e{+05}
@@
Additionally, we made computations and included also in Table 10 the eigenvalues of stable open-loop hovering [46]; the reader is encouraged to recall relevant comments/discussions about open-loop stability provided in Subsection III.A. To obtain the optimal–PID eigenvalues, the averaged model was first derived exactly as in the open-loop case. Classical linear-control theory was then applied to incorporate the PID loop and evaluate the closed-loop eigenvalues, following the procedure in [59]. All computational steps are available in the accompanying code repository in [60]. It is important to mention that there is a difference in the nature of the stability between PID/ES on one hand and the open-loop case on the other. In the open-loop configuration, the equilibrium need not correspond to hovering ($w=0$). Stability can occur at a non-zero $w$, leading to steady ascent or descent, as discussed in Section III.B; only by selecting the exact trim (numerically demanding) does one obtain open-loop hovering. By contrast, both the PID loop and the ES controller always stabilize to w=0w\!=\!0, so any disturbance is rejected and altitude is recovered. It can also be noted that in some instances the optimal PID closed-loop is less effective than the open-loop system—in the hawkmoth case, for example, as shown in Table 10 (but this should be put in the context that PID stabilization is hovering stabilization, which is not the case in the open-loop case. Overall, the eigenvalue results confirm the improvement in hovering stability by the natural hovering ES system using different objective functions compared to the open-loop case and optimal PID case. The results confirm that via extremum seeking control (ESC), hovering exhibits eigenvalues with more negative real components (i.e., shifts the eigenvalues further into the left-half complex plane), indicating stronger stability properties and more robustness.